### v1

@@ 7 Conclusion
Kernel density machines (KDM) is a comprehensive data-driven framework for estimating the density ratio (Radon-Nikodym derivative) $\frac{d\mathbb{Q}}{d\mathbb{P}}$ from i.i.d. samples of $\mathbb{P}$ and $\mathbb{Q}$ in a reproducing kernel Hilbert hypothesis space. KDM is computable in particular with large data sets, and allows both finite-sample and asymptotic inference. Accordingly, we illustrate its use within hypothesis testing and conditional distribution estimation. Given the reliance of learning problems on the data-generating probability law and the law induced by a model, KDM can be used in many scenarios ranging from transfer learning problems, and generative models to causal inference.

### v2

@@ 7 Conclusion
Kernel density machines (KDM) is a comprehensive data-driven framework for estimating the Radon–Nikodym derivative (density) $\frac{d\mathbb{Q}}{d\mathbb{P}}$ from i.i.d. samples of $\mathbb{P}$ and $\mathbb{Q}$ in a reproducing kernel Hilbert hypothesis space. KDM is computable in particular with large data sets, and allows both finite-sample and asymptotic inference. Accordingly, we illustrate its use within hypothesis testing and conditional distribution estimation. Given the reliance of learning problems on the data-generating probability law and the law induced by a model, KDM can be used in many scenarios ranging from transfer learning problems, and generative models to causal inference.